# On the Training Convergence of Transformers for In-Context Classification

### v1

@@ D.2 Notations
In this section, we introduce the following notations to analyze (7). We denote $\mu=\mu_{\tau}$, $\mu_{1}=\mu_{\tau,1}$, $\mu_{0}=\mu_{\tau,0}$ and $q=x_{\tau,\mathsf{query}}$. Then with probability $\mathbb{P}\left(y_{\tau,\mathsf{query}}=1\right)=1/2$ we have $q=\mu_{1}+v$, and with probability $\mathbb{P}\left(y_{\tau,\mathsf{query}}=0\right)=1/2$ we have $q=\mu_{0}+v$, where $v\sim\mathsf{N}(0,\Lambda)$. We define $p=\frac{2}{N}\sum_{i=1}^{N}y_{\tau,i}x_{\tau,i}$. Since with probability $\mathbb{P}\left(y_{\tau,i}=1\right)=1/2$ we have $x_{\tau,i}=\mu_{1}+v_{i}$, and with probability $\mathbb{P}\left(y_{\tau,i}=0\right)=1/2$ we have $x_{\tau,i}=\mu_{0}+v_{i}$, where $v_{i}\sim\mathsf{N}(0,\Lambda)$, we known $p=2N_{1}\mu_{1}/N-2N_{0}\mu_{0}/N+g$, where $g=\frac{2}{N}\sum_{i=1}^{N}v_{i}$, $g\sim\mathsf{N}(0,4\Lambda/N)$, $N_{1}\sim\mathsf{Bin}(N,1/2)$. Defining $h=N_{1}/N-1/2$, $u=2(\mu_{1}+\mu_{0})$, we have $N_{0}/N=1/2-h$ and
@@ D.3 Lemmas
Proof   Since $N_{1}\sim\mathsf{Bin}(N,1/2)$, the moment-generating function of $N_{1}$ is
@@
Proof   We vectorize $W$ as $\text{Vec}(W)\in\mathbb{R}^{d^{2}}$, where $\text{Vec}(W)_{i}=W_{t_{1}(i),t_{2}(i)}$, $t_{1}(x)=\lfloor(x-1)/d\rfloor+1,t_{2}(x)=((x-1)\text{ mod }d)+1$. Then, we have
@@
Similarly, we can prove the same conclusions for $\widetilde{L}(W)$. $\blacksquare$
Lemma D.3
@@
Proof   Let $a=\mu^{\top}\Lambda^{-1}q$, $s=\mu^{\top}Wq/2$, $r=(hu+g)^{\top}Wq/2$. Performing the Taylor expansion on (7), we have
@@
$=$ $\mathbb{E}\left[-\sigma(a)\log(\sigma(s))-(1-\sigma(a))\log(1-\sigma(s))\right]-\mathbb{E}\left[\left(\sigma(a)(1-\sigma(s))-(1-\sigma(a))\sigma(s))\right)r\right]$
@@
$=$ $\widetilde{L}(W)-\mathbb{E}\left[\left(\sigma(a)(1-\sigma(s))-(1-\sigma(a))\sigma(s))\right)r\right]+\mathbb{E}\left[\sigma(\xi(s,r))(1-\sigma(\xi(s,r)))r^{2}/2\right],$
@@
$\mathbb{E}\left[\sigma(\xi(s,r))(1-\sigma(\xi(s,r))r^{2}/2\right]$ $\leq\mathbb{E}\left[r^{2}\right]$
@@
Thus, when $N\to\infty$, we have $W^{*}\to 2\Lambda^{-1}$. Denoting $W^{*}=2(\Lambda^{-1}+G)$, we have $\|G\|_{\max}=O(1/\sqrt{N})$. $\blacksquare$
Lemma D.4
@@
Proof   According to Lemma D.2, the loss function $L(W)$ has a unique global minimizer $W^{*}$. We have
@@
$\mathbb{E}_{g,h}[bpq^{\top}]$ $=\mathbb{E}[\mu^{\top}Gq\mu q^{\top}+g^{\top}\Lambda^{-1}qgq^{\top}+g^{\top}Gqgq^{\top}+h^{2}u^{\top}Gquq^{\top}+h^{2}u^{\top}\Lambda^{-1}quq^{\top}]$
@@
Each term in $(i)$ contains two $G$’s. Thus, their max norms are at most $O(\|G\|_{\max}^{2})$. For each term in $(ii)$, it contains one $G$ and $h^{2}$ or contains one $G$ and two $g$. According to $\mathbb{E}[h^{2}]=1/(4N)$ in Lemma D.1, the max norm of terms with one $G$ and $h^{2}$ are smaller than $O(\|G\|_{\max}/N)$. Defining $\bar{g}=N^{1/2}\Lambda^{-1/2}g/2$, we have $\bar{g}\sim\mathsf{N}(0,I_{d})$ and $g=2N^{-1/2}\Lambda^{1/2}\bar{g}$. Thus, converting two $g$ to $\bar{g}$, we have a coefficient of $N^{-1}$. Therefore, the max norms of terms with one $G$ and two $g$ are also smaller than $O(\|G\|_{\max}/N)$. Therefore, for terms $(i),(ii)$, we have
@@
$\|\mathbb{E}[\sigma^{\prime\prime\prime}(\xi(a,b))b^{3}pq^{\top}/3!]\|_{\max}\leq\max_{z\in\mathbb{R}}|\sigma^{\prime\prime\prime}(z)|/3!\cdot\max_{ij}\mathbb{E}\left[\left|b^{3}p_{i}q_{j}\right|\right]$
@@
$\|\mathbb{E}[\sigma^{\prime\prime\prime}(\xi(a,b))b^{3}pq^{\top}/3!]\|_{\max}$ $\leq\max_{ij}\mathbb{E}\left[\left|b^{3}p_{i}q_{j}\right|\right]\cdot\max_{z}|\sigma^{\prime\prime\prime}(z)|/3!$
@@
$\|G\|_{\max}\leq\frac{1}{N}\|S^{-1}\left(\mathbb{E}[\sigma^{\prime}(a)(4qq^{\top}+uu^{\top}\Lambda^{-1}qq^{\top}/4)+\sigma^{\prime\prime}(a)((u^{\top}\Lambda^{-1}q)^{2}\mu q^{\top}/8+2q^{\top}\Lambda^{-1}q\mu q^{\top})]\right)\|_{\max}+o(1/N).$
$\blacksquare$
Lemma D.5
@@
Proof   The Hessian matrix of the loss function is
@@
where $(a)$ is due to the Cauchy–Schwarz inequality. Thus, $\nabla^{2}L(W)\preceq lI_{d}$ and $L(W)$ is $l$-smooth, where $l$ is a constant smaller than $\frac{1}{4}\sum_{i\in[d^{2}]}\mathbb{E}[(p_{t_{1}(i)}q_{t_{2}(i)})^{2}]$. $\blacksquare$
@@
Proof   According to Lemma D.4, the global minimizer of $L(W)$ is $W^{*}=2(\Lambda^{-1}+G)$, where
@@
where $\kappa=l/\alpha$. $\blacksquare$
@@ E.1 Notations
Proof   The output of the trained transformer is
@@
$\blacksquare$
Remark E.1
We note that Theorem 3.2 requires Assumption 3.2 to hold. For example, we need the covariance $\Lambda$ in training and testing to be the same. A similar consistency requirement of the covariance $\Lambda$ in training and testing had also been observed for in-context linear regression in Zhang et al. [2023a].
Here, we discuss the consequences when Assumption 3.2 does not hold. For example, suppose the labels of our data in test prompts are not balanced where $\mathbb{P}\left(y=1\right)=p_{1},\mathbb{P}\left(y=-1\right)=p_{0}$. Besides, $\mu_{0},\mu_{1}$ do not have the same $\Lambda^{-1}$ weighted norm, and the covariance matrix of test data is $\Gamma\neq\Lambda$. Then, as $N,M\to\infty$, we have
$\frac{2}{M}\sum_{i=1}^{M}y_{i}x_{i}^{\top}\to 2(p_{1}\mu_{1}-p_{0}\mu_{0})^{\top},$
and
On the other hand, the distribution of the ground truth label is
$\mathbb{P}\left(y_{\mathsf{query}}=1\right)=\sigma((\mu_{1}-\mu_{0})^{\top}\Gamma^{-1}x_{\mathsf{query}}+(\mu_{1}^{\top}\Lambda^{-1}\mu_{1}-\mu_{0}^{\top}\Lambda^{-1}\mu_{0})/2+\log(p_{1}/p_{0})).$
Define $z\triangleq(\mu_{1}-\mu_{0})^{\top}\Gamma^{-1}x_{\mathsf{query}}+(\mu_{1}^{\top}\Lambda^{-1}\mu_{1}-\mu_{0}^{\top}\Lambda^{-1}\mu_{0})/2+\log(p_{1}/p_{0})$ and $\hat{z}\triangleq 2(p_{1}\mu_{1}-p_{0}\mu_{0})^{\top}\Lambda^{-1}x_{\mathsf{query}}$. Then, we can notice that unless $\hat{z}=z$ or $\left|\sigma(\hat{z})-\sigma(z)\right|$ is sufficiently small, the transformer cannot correctly perform the in-context binary classification.
@@ F.2 Notations
Proof   Since $(N_{1},N_{2},\dots,N_{c})\sim\mathsf{Multin}(N,1/c)$, the moment-generating function of $(N_{1},N_{2},\dots,N_{c})$ is
@@
Proof   Since $g_{k}\sim\mathsf{N}(0,c^{2}N_{k}\Lambda/N^{2})$ and $\bar{g}_{k}\sim\mathsf{N}(0,c^{2}N_{k}I_{d}/N^{2})=\mathsf{N}(0,(c/N+c^{2}h_{k}/N)I_{d})$, we have
@@
Proof   We vectorize $W$ as $\text{Vec}(W)\in\mathbb{R}^{d^{2}}$, where $\text{Vec}(W)_{i}=W_{t_{1}(i),t_{2}(i)}$, $t_{1}(x)=\lfloor(x-1)/d\rfloor+1,t_{2}(x)=((x-1)\text{ mod }d)+1$. Then, we have
@@
Similarly, we can prove the same conclusions for $\widetilde{L}(W)$. $\blacksquare$
Lemma F.4
@@
Proof   Let $a=\mu^{\top}\Lambda^{-1}q$, $s=\mu^{\top}Wq/c$, $r=(\mu_{h}+g)^{\top}Wq/c$, $a_{k}=\mu_{k}^{\top}\Lambda^{-1}q$, $s_{k}=\mu_{k}^{\top}Wq/c$, $r_{k}=(ch_{k}\mu_{k}+g_{k})^{\top}Wq/c$. Performing the Taylor expansion on (16), we have
@@
$\left|\widetilde{L}(W)-L(W)\right|\leq$ $c\sum_{l=1}^{c}\mathbb{E}\left[|r_{l}|\right]$
@@
$\nabla\widetilde{L}(W)=\mathbb{E}\left[\sum_{k=1}^{c}-\mathrm{softmax}(\mu^{\top}\Lambda^{-1}q)_{k}\log(\mathrm{softmax}(\mu^{\top}Wq/c)_{k})\right]=0$
@@
Thus, when $N\to\infty$, we have $W^{*}\to c\Lambda^{-1}$. Denoting $W^{*}=c(\Lambda^{-1}+G)$, we have $\|G\|_{\max}=O(N^{-1/4})$. $\blacksquare$
Lemma F.5
@@
Proof   According to Lemma F.3, the loss function $L(W)$ has a unique global minimizer $W^{*}$. We have
@@
Ignoring constants other than $c,N$, we have $\|G\|_{\max}\leq O(c/N)$. $\blacksquare$
Lemma F.6
@@
Proof   The Hessian matrix of the loss function is
@@
$z^{\top}\nabla^{2}L(W)z=$ $\mathbb{E}\left[\frac{1}{c^{2}}\sum_{k=2}^{c}\sum_{l=1}^{k-1}\mathrm{softmax}(P^{\top}Wq/c)_{k}\mathrm{softmax}(P^{\top}Wq/c)_{l}\left(\sum_{a\in[d^{2}]}z_{a}(p_{k}-p_{l})_{t_{1}(a)}q_{t_{2}(a)}\right)^{2}\right]$
@@
where $(a)$ is due to the Cauchy–Schwarz inequality. Thus, $\nabla^{2}L(W)\preceq lI_{d}$ and $L(W)$ is $l$-smooth, where $l$ is a constant smaller than $\frac{1}{c^{2}}\sum_{k=2}^{c}\sum_{l=1}^{k-1}\sum_{i\in[d^{2}]}\mathbb{E}[((p_{k}-p_{l})_{t_{1}(i)}q_{t_{2}(i)})^{2}]$. $\blacksquare$
Theorem F.1 (Formal statement of Theorem 4.1)
@@
Proof   According to Lemma F.5, the global minimizer of $L(W)$ is $W^{*}=c(\Lambda^{-1}+G)$, where
@@
After $T\geq 2\kappa\log(N\cdot\|W^{0}-W^{*}\|_{F})$ gradient steps, we have $\widehat{W}=W^{T}=c(\Lambda^{-1}+G+H^{T}/c)=2(\Lambda^{-1}+\widehat{G})$, where $\widehat{G}=G+H^{T}/c$, $\|H^{T}\|_{\max}\leq\exp(-T/\kappa)\cdot\|W^{0}-W^{*}\|_{F}^{2}\leq 1/N$. Thus, $\|\widehat{G}\|_{\max}\leq\|G\|_{\max}+\|H^{T}\|_{\max}=O(c/N)$. $\blacksquare$

### v3

@@ D.2 Notations
In this section, we introduce the following notations to analyze (7). We denote $\mu=\mu_{\tau,1}-\mu_{\tau,0}$, $\mu_{1}=\mu_{\tau,1}$, $\mu_{0}=\mu_{\tau,0}$ and $q=x_{\tau,\mathsf{query}}$. Then with probability $\mathbb{P}\left(y_{\tau,\mathsf{query}}=1\right)=1/2$ we have $q=\mu_{1}+v$, and with probability $\mathbb{P}\left(y_{\tau,\mathsf{query}}=0\right)=1/2$ we have $q=\mu_{0}+v$, where $v\sim\mathsf{N}(0,\Lambda)$. We define $p=\frac{2}{N}\sum_{i=1}^{N}y_{\tau,i}x_{\tau,i}$. Since with probability $\mathbb{P}\left(y_{\tau,i}=1\right)=1/2$ we have $x_{\tau,i}=\mu_{1}+v_{i}$, and with probability $\mathbb{P}\left(y_{\tau,i}=0\right)=1/2$ we have $x_{\tau,i}=\mu_{0}+v_{i}$, where $v_{i}\sim\mathsf{N}(0,\Lambda)$, we known $p=2N_{1}\mu_{1}/N-2N_{0}\mu_{0}/N+g$, where $g=\frac{2}{N}\sum_{i=1}^{N}v_{i}$, $g\sim\mathsf{N}(0,4\Lambda/N)$, $N_{1}\sim\mathsf{Bin}(N,1/2)$. Defining $h=N_{1}/N-1/2$, $u=2(\mu_{1}+\mu_{0})$, we have $N_{0}/N=1/2-h$ and
@@ D.3 Lemmas
Proof.
Since $N_{1}\sim\mathsf{Bin}(N,1/2)$, the moment-generating function of $N_{1}$ is
@@
Proof.
We vectorize $W$ as $\text{Vec}(W)\in\mathbb{R}^{d^{2}}$, where $\text{Vec}(W)_{i}=W_{t_{1}(i),t_{2}(i)}$, $t_{1}(x)=\lfloor(x-1)/d\rfloor+1,t_{2}(x)=((x-1)\text{ mod }d)+1$. Then, we have
@@
Similarly, we can prove the same conclusions for $\widetilde{L}(W)$. ∎
Lemma D.3.
@@
Proof.
Let $a=\mu^{\top}\Lambda^{-1}q$, $s=\mu^{\top}Wq/2$, $r=(hu+g)^{\top}Wq/2$. Performing the Taylor expansion on (7), we have
@@
$=$ $\mathbb{E}\left[-\sigma(a)\log(\sigma(s))-(1-\sigma(a))\log(1-\sigma(s))\right]$
$-\mathbb{E}\left[\left(\sigma(a)(1-\sigma(s))-(1-\sigma(a))\sigma(s))\right)r\right]$
@@
$=$ $\widetilde{L}(W)-\mathbb{E}\left[\left(\sigma(a)(1-\sigma(s))-(1-\sigma(a))\sigma(s))\right)r\right]$
$+\mathbb{E}\left[\sigma(\xi(s,r))(1-\sigma(\xi(s,r)))r^{2}/2\right],$
@@
$\mathbb{E}\left[\sigma(\xi(s,r))(1-\sigma(\xi(s,r))r^{2}/2\right]$
$\leq$ $\mathbb{E}\left[r^{2}\right]$
@@
Thus, when $N\to\infty$, we have $W^{*}\to 2\Lambda^{-1}$. Denoting $W^{*}=2(\Lambda^{-1}+G)$, we have $\|G\|_{\max}=O(1/\sqrt{N})$. ∎
Lemma D.4.
@@
Proof.
According to Lemma D.2, the loss function $L(W)$ has a unique global minimizer $W^{*}$. We have
@@
$\mathbb{E}_{g,h}[bpq^{\top}]$
$=$ $\mathbb{E}[\mu^{\top}Gq\mu q^{\top}+g^{\top}\Lambda^{-1}qgq^{\top}+g^{\top}Gqgq^{\top}+h^{2}u^{\top}Gquq^{\top}+h^{2}u^{\top}\Lambda^{-1}quq^{\top}]$
@@
For each term in $(i)$, it contains two $G$. Thus, their max norms are at most smaller than $O(\|G\|_{\max}^{2})$. For each term in $(ii)$, it contains one $G$ and $h^{2}$ or contains one $G$ and two $g$. According to $\mathbb{E}[h^{2}]=1/(4N)$ in Lemma D.1, the max norm of terms with one $G$ and $h^{2}$ are smaller than $O(\|G\|_{\max}/N)$. Defining $\bar{g}=N^{1/2}\Lambda^{-1/2}g/2$, we have $\bar{g}\sim\mathsf{N}(0,I_{d})$ and $g=2N^{-1/2}\Lambda^{1/2}\bar{g}$. Thus, converting two $g$ to $\bar{g}$, we have a coefficient of $N^{-1}$. Therefore, the max norms of terms with one $G$ and two $g$ are also smaller than $O(\|G\|_{\max}/N)$. Therefore, for terms $(i),(ii)$, we have
@@
$\|\mathbb{E}[\sigma^{\prime\prime\prime}(\xi(a,b))b^{3}pq^{\top}/3!]\|_{\max}$
$\leq$ $\max_{z\in\mathbb{R}}|\sigma^{\prime\prime\prime}(z)|/3!\cdot\max_{ij}\mathbb{E}\left[\left|b^{3}p_{i}q_{j}\right|\right]$
@@
$\|\mathbb{E}[\sigma^{\prime\prime\prime}(\xi(a,b))b^{3}pq^{\top}/3!]\|_{\max}$
$\leq$ $\max_{ij}\mathbb{E}\left[\left|b^{3}p_{i}q_{j}\right|\right]\cdot\max_{z}|\sigma^{\prime\prime\prime}(z)|/3!$
@@
$\|G\|_{\max}$
$\leq$ $\frac{1}{N}\|S^{-1}\left(\mathbb{E}[\sigma^{\prime}(a)(4qq^{\top}+uu^{\top}\Lambda^{-1}qq^{\top}/4)+\sigma^{\prime\prime}(a)((u^{\top}\Lambda^{-1}q)^{2}\mu q^{\top}/8+2q^{\top}\Lambda^{-1}q\mu q^{\top})]\right)\|_{\max}$
$+o(1/N).$
Lemma D.5.
@@
Proof.
The Hessian matrix of the loss function is
@@
where $(a)$ is due to the Cauchy–Schwarz inequality. Thus, $\nabla^{2}L(W)\preceq lI_{d}$ and $L(W)$ is $l$-smooth, where $l$ is a constant smaller than $\frac{1}{4}\sum_{i\in[d^{2}]}\mathbb{E}[(p_{t_{1}(i)}q_{t_{2}(i)})^{2}]$. ∎
@@
Proof.
According to Lemma D.4, the global minimizer of $L(W)$ is $W^{*}=2(\Lambda^{-1}+G)$, where
@@
where $\kappa=l/\alpha$. ∎
@@ E.1 Notations
Proof.
The output of the trained transformer is
@@ F.2 Notations
Proof.
Since $(N_{1},N_{2},\dots,N_{c})\sim\mathsf{Multin}(N,1/c)$, the moment-generating function of $(N_{1},N_{2},\dots,N_{c})$ is
@@
Proof.
Since $g_{k}\sim\mathsf{N}(0,c^{2}N_{k}\Lambda/N^{2})$ and $\bar{g}_{k}\sim\mathsf{N}(0,c^{2}N_{k}I_{d}/N^{2})=\mathsf{N}(0,(c/N+c^{2}h_{k}/N)I_{d})$, we have
@@
Proof.
We vectorize $W$ as $\text{Vec}(W)\in\mathbb{R}^{d^{2}}$, where $\text{Vec}(W)_{i}=W_{t_{1}(i),t_{2}(i)}$, $t_{1}(x)=\lfloor(x-1)/d\rfloor+1,t_{2}(x)=((x-1)\text{ mod }d)+1$. Then, we have
@@
Similarly, we can prove the same conclusions for $\widetilde{L}(W)$. ∎
Lemma F.4.
@@
Proof.
Let $a=\mu^{\top}\Lambda^{-1}q$, $s=\mu^{\top}Wq/c$, $r=(\mu_{h}+g)^{\top}Wq/c$, $a_{k}=\mu_{k}^{\top}\Lambda^{-1}q$, $s_{k}=\mu_{k}^{\top}Wq/c$, $r_{k}=(ch_{k}\mu_{k}+g_{k})^{\top}Wq/c$. Performing the Taylor expansion on (16), we have
@@
$\left|\widetilde{L}(W)-L(W)\right|$
$\leq$ $c\sum_{l=1}^{c}\mathbb{E}\left[|r_{l}|\right]$
@@
$\nabla\widetilde{L}(W)=\mathbb{E}[\sum_{k=1}^{c}-\mathrm{softmax}(\mu^{\top}\Lambda^{-1}q)_{k}\log(\mathrm{softmax}(\mu^{\top}Wq/c)_{k})]=0$
@@
Thus, when $N\to\infty$, we have $W^{*}\to c\Lambda^{-1}$. Denoting $W^{*}=c(\Lambda^{-1}+G)$, we have $\|G\|_{\max}=O(N^{-1/4})$. ∎
Lemma F.5.
@@
Proof.
According to Lemma F.3, the loss function $L(W)$ has a unique global minimizer $W^{*}$. We have
@@
$\|G\|_{\max}$
$\leq$ $\frac{1}{N}\Bigg{\|}S^{-1}\mathbb{E}\Bigg{[}\sum_{k,l=1}^{c}\frac{\partial\zeta_{k}(a)}{\partial a_{l}}(c\delta_{kl}-1)\mu_{k}\mu_{l}^{\top}\Lambda^{-1}qq^{\top}+\sum_{k=1}^{c}\frac{\partial\zeta_{k}(a)}{\partial a_{k}}cqq^{\top}$
@@
Ignoring constants other than $c,N$, we have $\|G\|_{\max}\leq O(c/N)$. ∎
Lemma F.6.
@@
Proof.
The Hessian matrix of the loss function is
@@
$z^{\top}\nabla^{2}L(W)z$
@@
where $(a)$ is due to the Cauchy–Schwarz inequality. Thus, $\nabla^{2}L(W)\preceq lI_{d}$ and $L(W)$ is $l$-smooth, where $l$ is a constant smaller than $\frac{1}{c^{2}}\sum_{k=2}^{c}\sum_{l=1}^{k-1}\sum_{i\in[d^{2}]}\mathbb{E}[((p_{k}-p_{l})_{t_{1}(i)}q_{t_{2}(i)})^{2}]$. ∎
Theorem F.7 (Formal statement of Theorem 4.3).
@@
Proof.
According to Lemma F.5, the global minimizer of $L(W)$ is $W^{*}=c(\Lambda^{-1}+G)$, where
$\|G\|_{\max}$
$\leq$ $\frac{1}{N}\Bigg{\|}S^{-1}\mathbb{E}\Bigg{[}\sum_{k,l=1}^{c}\frac{\partial\zeta_{k}(a)}{\partial a_{l}}(c\delta_{kl}-1)\mu_{k}\mu_{l}^{\top}\Lambda^{-1}qq^{\top}+\sum_{k=1}^{c}\frac{\partial\zeta_{k}(a)}{\partial a_{k}}cqq^{\top}$
@@
After $T\geq 2\kappa\log(N\cdot\|W^{0}-W^{*}\|_{F})$ gradient steps, we have $\widehat{W}=W^{T}=c(\Lambda^{-1}+G+H^{T}/c)=2(\Lambda^{-1}+\widehat{G})$, where $\widehat{G}=G+H^{T}/c$, $\|H^{T}\|_{\max}\leq\exp(-T/\kappa)\cdot\|W^{0}-W^{*}\|_{F}^{2}\leq 1/N$. Thus, $\|\widehat{G}\|_{\max}\leq\|G\|_{\max}+\|H^{T}\|_{\max}=O(c/N)$. ∎